OpenStudy (ria23):

The circle given by x^2+y^2-2y-11=0 can be written in standard form like this: x^2+(y-k)^2=12. What is the value of k in this equation? I'm not really sure what the question is asking me...

OpenStudy (anonymous):

this is the cirlcle's formula/equation .. (x-h)+(y-k)^2=r^2

OpenStudy (ria23):

So, I'm not sure if I explain what k stands for or what I'm supposed to plug in for k... Assuming that it means what II plug in because of "What is the value of k in this equation?" Would my k value be -2?

OpenStudy (anonymous):

let me try to solve it

OpenStudy (ria23):

Alright. No problem. n_n Thank yhu for helping me~

OpenStudy (anonymous):

what is the subject title?

OpenStudy (ria23):

Pre-Calculus

OpenStudy (anonymous):

x^2+y^2-2y-11=0 x^2+y^2-2y=-11 x^2+y^2-2y-1=-11 x^2+y^2-2y-1=-11-1 x^2+(y-1)^2=-12 from this equation(x-h)=(y-k)=r^2 the value of y would be -1 .. I dont know if you're using the same approach with mine.

OpenStudy (ria23):

Haha, I guess I wasn't. >~< Thank yhu so much for the help. And taking the time to work that out

OpenStudy (anonymous):

haha.. Im using analytic geometry for this one.

OpenStudy (ria23):

That's actually how my teacher worked it out. I can see where I went wrong by looking at what yhu did. I missed subtracting the 1

OpenStudy (ria23):

Adding the 1* My bad.

OpenStudy (anonymous):

so,what value did you got?

OpenStudy (ria23):

I got positive 2 to start with, I forgot to add the 1 When I did it again I got -1

OpenStudy (ria23):

Yhu had an error tho. \[x^{2}+y^{2}-2y-11=0\]\[x^{2}+y^{2}-2y=11\]\[x^{2}+y^{2}-2y+1=11+1\]\[x^{2}+(y-1)^{2}=12\]

OpenStudy (anonymous):

what part?

OpenStudy (ria23):

Yhu want to add 11 to each side, which will give yhu positive 11 on the other side of the = sign.